Titmouse lab can have on average 200 lab tests done per day. There are 3 working shifts, each is 6 hours. The lab has 15 workers for each shift, and pays each worker $20 per hour. The utility cost for the lab is $50 per hour. No utility cost will occur when the lab is not in operation. Please calculate the lab's labor cost per day.
Answer #1

Labor cost per shift per employee = Labor cost per hour * hours in a shift

Labor cost per shift per employee = 20 * 6 = $120

Labor cost per shift = Number of Employees in each shift * Labor cost per shift per employee

Labor cost per shift = 15 * 120 = $1800

Labor cost per day = Shifts per day * Labor cost per shift

Labor cost per day = 3* 1800

Labor cost per day = $ 5400